OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (-2, -3) and is parallel to the line y = x + 9?

OpenStudy (anonymous):

Because the new line is parallel, it will have the same slope. You can determine the slope of the first line by inspecting it as being in the form y = mx + b where m is the slope. Then for a point \[(x _{1}, y _{1})\] which you are given, use point-slope of \[y - y _{1} = m(x - x _{1})\]

OpenStudy (anonymous):

As for standard form for the equation of the line, once you solve the point-slope form, rearrange to put the equation in ax + by = c.

OpenStudy (anonymous):

i dont understand....

OpenStudy (anonymous):

Which part?

OpenStudy (anonymous):

all of it. im terrible at math

OpenStudy (anonymous):

np. Your original line is y = x + 9. Recognizing that this equation is in slope-intercept form which y = mx + b, can you tell what the value of "m" is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I'll reqrite the equation of the line in a slightly different form. y = (1)x + 9. Comparing that to y = mx + b, can you see what "m" is now?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

Yes! Very good. Because you needed the equation rewritten, before we go on, I need to remind you that multiplication by "1" preserves the number. As in x = (1)x. It's important to get that concept down before going on. All good?

OpenStudy (anonymous):

Yes? No?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Good. You are given a point (-2, -3) which is in the form of \[(x _{1}, y _{1})\] which nothing more than this is a specific point with a specific value for x and y. That's why there are subscripts. They represent a specific set value for x and y. So, now, substitute those values into the equation form my first post. Write it out so I know your are following all of this.

OpenStudy (anonymous):

i feel like an idiot

OpenStudy (anonymous):

np, you're not. Everyone starts at the bottom with math and works his way up. We all have to or have already gone through the same steps. They can be hard at first, but I'll help as much as I can.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

We'll take it slow. np. Let's just look at the left-hand side for now which is \[y - y _{1}\] You are given the point (-2, -3), which corresponds to the form \[(x _{1}, y _{1})\] So, x1 is -2 and y1 is -3. So, with that information, are you able to substitute the y1 value into the expression\[y - y _{1}\] ?

OpenStudy (anonymous):

uh i think so

OpenStudy (anonymous):

Ok, good. Write that out so you get a feel for this.

OpenStudy (anonymous):

ah i give up :/

OpenStudy (anonymous):

Can't you put -3 into \[y - y _{1}\] ?

OpenStudy (anonymous):

I'll give you a hint: the expression becomes y - (-3). Can you simplify that a little?